OpenStudy (anonymous):

if 2-cos^2 θ = 3sinθcosθ, sinθ not equal to cosθ, then tanθ is ? 1) 1/2 2) 0 3) 2/3 4) 1/3

OpenStudy (anonymous):

ok so lets try to change this "2" into "2sin^2 + 2cos^2"

OpenStudy (anonymous):

and divide both sides of the equation by cos^2 what do we get ?

OpenStudy (anonymous):

are you there ?

OpenStudy (anonymous):

2 tan^2θ +1= 3tanθ @coolsector

OpenStudy (shubhamsrg):

not 2 tan^2 @,,you get 2 sec^2 @ ..right?

OpenStudy (shubhamsrg):

and sec^2 @ = 1 + tan^2 @ use this and solve for tan @

OpenStudy (anonymous):

\[2-\cos^2 θ = 3\sinθ\cosθ\]divide by \(\cos^2 \theta\)\[\frac{2}{\cos^2 θ}-1 = 3\frac{\sinθ\cosθ}{\cos^2 \theta}\]\[2(1+\tan^2 \theta)-1=3\tan \theta\]\[2\tan^2 \theta-3\tan \theta+1=0\]a quadratic in terms of \(\tan \theta\)

OpenStudy (aravindg):

oops i am late

OpenStudy (anonymous):

dhiwagm you got the same equation as mukushla got .. this is good.. he solved it completely anyway so nothing much for you to do